OpenStudy (anonymous):

what are the formulas for the following.. 1. pyramid 2. cylinder 3. prism 4. sphere 5. cone

OpenStudy (anonymous):

give me one second..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Prisms Volume = Base X Height v=bh Surface = 2b + Ph (b is the area of the base P is the perimeter of the base) Cylinder Volume = pr2 x height V = pr2 h Surface = 2p radius x height S = 2prh + 2pr2 Pyramid V = 1/3 bh b is the area of the base Surface Area: Add the area of the base to the sum of the areas of all of the triangular faces. The areas of the triangular faces will have different formulas for different shaped bases. Cones Volume = 1/3 pr2 x height V= 1/3 pr2h Surface = pr2 + prs S = pr2 + prs =pr2 + pr Sphere Volume = 4/3 pr3 V = 4/3 pr3 Surface = 4pr2 S = 4pr2

OpenStudy (anonymous):

did that help? can I have a medal?

OpenStudy (anonymous):

Thanks hope that really helped!

OpenStudy (anonymous):

your welcome and it didnt really help the answer choices are L.A.=pi.r.l, V=3/4pi.r^3, T.A=2pi.r.h.+2pi.r^2, V=1/3B.h, L.A.=p.h

OpenStudy (anonymous):

so you where looking for something else then..oh.. ok

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

sorry